Kevin’s room is 12ft. x 15ft. Kevin draws a scale model of his room. If the width of his drawing is 2 inches, how long should he make the length?
Answer by rajagopalan(174)

Kevin's Room Size=12 by 15
Scaled drawing width=2
Let Lenth=x
Then....W:L=w:x
12:15 =2:x
product of extreme numbers=product of Mid numbers
12x=15*2
12x=30
x=30/12=2.5 inches
The length should be 2.5 inches
Ok